# Variable Neighborhood Searching Rerandomization

### v1

@@ 5.2 Structured Randomized Experiments
This section considers the three types of structured experiments described in Section 4. We simulate the covariates and potential outcomes in the same way as in Section 5.1, except that we set $Y_{i}(1)=Y_{i}(0)+0.15\sqrt{Var[Y_{i}(0)]}$ under the alternative to ensure the power of different methods remains within a moderate range. For sequential experiments, we set the number of groups as $K=2$ and consider two group sizes, $n_{k}\in\{100,500\}$. We use the strategy proposed by Zhou et al. (2018) to determine $a_{k}$. Specifically, we set $(s_{1},s_{2})=(239,761)$ when $p=50$ and $(s_{1},s_{2})=(264,736)$ when $p=250$, where $a_{k}=\frac{n_{k}}{n_{[k]}}q_{k}$ and $q_{k}$ is the $\frac{1}{s_{k}}$ quantile of a non-central chi-square distribution with $p$ degrees of freedom and a non-central parameter $\frac{n_{[k-1]}}{n_{k}}M_{[k-1]}$. For stratified experiments, we set the stratum number as $K=2$ and consider two stratum sizes, $n_{k}\in\{100,500\}$. For cluster randomized experiments, we set the cluster size as $n_{k}=2$ and consider two options for the number of clusters, $K\in\{100,500\}$. We set $p_{a}=P(\chi_{p}^{2}<a)=10^{-3}$ for both stratified randomized experiments and cluster randomized experiments and set $p=n/4$ for all settings in this section.
Table 2 shows the results when $n=1000$. The complete results are provided in the Supplement. For clarity, we use a unified name to refer to different method variants. Since PSRR is not extended to stratified experiments and cluster experiments, it is excluded from the two schemes. The overall findings are consistent with the results in Section 5.1. VNSRR substantially reduces the computational burden of rerandomization, especially for high-dimensional datasets and in stratified randomized experiments and cluster randomized experiments, where ARSRR is the only method available to perform rerandomization. VNSRR maintains the statistical properties of rerandomization in all the settings and significantly improves statistical inference compared to CR.
Table 2: Statistical and computational performance of different methods in structured randomized experiments.
@@
Bias SD Size Power Coverage Length
@@
VNSRR 3.3 131 10.1 82 89.9 436 3.9
5.3 A Clinical Trail Example
In addition to the simulation studies, we analyze a real dataset from a phase 1 clinical trial to illustrate the advantages of VNSRR. This dataset was analyzed by Zhu and Liu (2022) as well. The goal of this trial was to evaluate the interactions between intravenous methamphetamine and oral reserpine. In this trial, 20 participants were randomly assigned to the treatment group, and 10 participants were assigned to the control group. We aim to balance better three covariates, age, weight, and pre-treatment heart rate, which are significantly unbalanced in the original study, to better infer the treatment effects on post-treatment heart rate.
We set $a$ as the 0.001 quantile of $\chi_{3}^{2}$ and generate 10,000 acceptable assignments using each method. RR takes approximately 2.3 minutes to complete the task, and PSRR takes more than half an hour. In contrast, VNSRR requires only 12 seconds, which demonstrates that VNSRR accelerates the rerandomization process significantly. Based on the original dataset, we simulate semi-synthetic data and examine the statistical performance of each method. The results are provided in the Supplement.
@@ 6 Conclusion
In this work, we propose a novel VNSRR method to perform rerandomization efficiently in various types of randomized experiments. We derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under VNSRR. Our empirical results demonstrate that VNSRR maintains the appealing statistical properties of rerandomization and significantly reduces the computational cost compared to existing methods. A future direction is to extend the proposed method to experiments with tiers of covariates. Multi-objective optimization techniques are required to search for assignments that simultaneously balance the covariates within each tier.
@@
In Appendix A, we provide the pseudocode of the three VNSRR variants. In Appendix B, we state and prove all the theoretical results. In Appendix C, we introduce the Fisher randomization test and randomization-based confidence intervals. In Appendix D, we show the simulation results not included in the main text.
Appendix A Algorithmic Description of VNSRR Variants
A.1 Algorithmic Description of SeqVNSRR
The pseudocode of SeqVNSRR is presented in Algorithm 2. As mentioned in the main text, SeqVNSRR performs VNSRR in each experiment stage sequentially while fixing the assignments obtained in previous stages.
A.2 Algorithmic Description of StratVNSRR
The pseudocode of StratVNSRR is presented in Algorithm 3. StratVNSRR differs from VNSRR in that there are linear constraints within each stratum, $\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk}$. Therefore, StratVNSRR only exchanges units in the same stratum to ensure that the aforementioned linear constraints are always satisfied. StratVNSRR does not loop over the strata sequentially. Instead, after selecting the pairs of units to examine from each stratum, StratVNSRR aggregates the lists of pairs and randomly permutes the combined list. In this way, we ensure that all strata are equally treated, allowing for flexible exploration of the neighborhood without a predetermined ordering of the strata or the units.
A.3 Algorithmic Description of ClustVNSRR
The pseudocode of ClustVNSRR is presented in Algorithm 4. ClustVNSRR treats each cluster as a whole and uses the clusters as the operational units in the search process. In other words, ClustVNSRR essentially conducts VNSRR at the cluster level rather than the unit level. Recall that we denote the assignment status of the clusters as a $K$-dimensional vector $U$ and all units’ assignment status as a $n$-dimensional vector $W$. We have the following constraint imposed by the cluster structure of the experiment. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit. Since $W$ is determined once we generate $U$, we denote $W(U)$ as the mapping from $U$ to $W$ in Algorithm 4.
@@ A.3 Algorithmic Description of ClustVNSRR
Algorithm 2 Sequential Variable neighborhood searching rerandomization
Input: Covariates $X$, index sets of units in each stratum $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of units to assign to the two arms in each stratum $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, threshold $a$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$, $L_{total}=\sum_{k}L_{k}$, $S_{total}=\sum_{k}S_{k}$.
@@
      Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{L_{total}})$ with random ordering;
       Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{L_{total}})$. The ordering of $(Z_{1},\ldots,Z_{L_{total}})$ is consistent with that of $(O_{1},\ldots,O_{L_{total}})$;
       Set $\widetilde{S}=S_{total}$;
       for $l=1,\ldots,L_{total}$ do
@@
            Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{S_{total}})$;
             Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{S_{total}})$;
             for $s=1,\ldots,S_{total}$ do
@@
Algorithm 3 Stratified Variable neighborhood searching rerandomization
@@
Algorithm 4 Cluster Variable neighborhood searching rerandomization
@@ Appendix B Additional Theoretical Results and Proofs
VNSRR, StratVNSRR, and ClustVNSRR all assign treatment in an offline manner. In contrast, SeqVNSRR performs online allocation and needs to be handled differently. We will first establish the theoretical properties of VNSRR, StratVNSRR, and ClustVNSRR, and then obtain the results of SeqVNSRR.
For the convenience of readers, we first restate the conditions we propose in the main text, which are related to the offline methods.
Condition 1
@@
Condition 4
$n_{tk}=n_{ck}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
Condition 5
@@
Property 1
$\mathcal{F}=1-\mathcal{F}$.
Property 2
For any $w\in\mathcal{F}$,
$P\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
Property 3
There exists a sequence of functions $\{u_{t}\}_{t\geq 1}$, such that for any $t$ and $(w^{(0)},\ldots,w^{(t)})$ that may appear consecutively in the generated chain,
$P\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
In this section, we first prove Theorem 1 and Theorem 2 based on the abovementioned conditions and properties.
Theorem 1
If $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then $E[\widehat{\tau}]=\tau$.
Theorem 2
Under Condition 1 and Condition 2, if $W$ is generated by a rerandomization algorithm satisfying Property 1, Property 2 and Property 3, then
$\frac{V_{CR}-Var[\widehat{\tau}]}{V_{CR}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where $R^{2}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}V_{CR}}$.
We then verify that VNSRR, StratVNSRR, and ClustVNSRR satisfy the requirements by proving the following propositions. Their unbiasedness and variance reduction are immediately evident as corollaries of Theorem 1 and Theorem 2.
Proposition 1
VNSRR satisfies Property 2 and Property 3. Under Condition 3, VNSRR satisfies Property 1.
Proposition 2
StratVNSRR satisfies Property 2 and Property 3. Under Condition 4, StratVNSRR satisfies Property 1.
Proposition 3
ClustVNSRR satisfies Property 2 and Property 3. Under Condition 5, ClustVNSRR satisfies Property 1.
We also show that the statistical properties of CR, ARSRR, and PSRR can be established in this way as well based on the following propositions.
Proposition 4
CR satisfies Property 1, Property 2 and Property 3.
Proposition 5
ARSRR satisfies Property 2 and Property 3. Under Condition 3, ARSRR satisfies Property 1.
Proposition 6
PSRR satisfies Property 2 and Property 3. Under Condition 3, PSRR satisfies Property 1.
Consequently, we will establish the statistical properties of SeqVNSRR by extending the conditions and properties we consider in the offline setting to the online setting.
@@ B.1 Proof of Theorem 1
$P\left(W^{(0)}=w^{(0)}\right)=P\left(W^{(0)}=1-w^{(0)}\right),$
@@
$\hskip 15.0ptP\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)$
@@
$=P\left(W^{(t)}=1-w^{(t)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(t-1)}=1-w^{(t-1)}\right),$
@@
$\hskip 15.0ptP\left(W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)},W^{(T)}=w^{(T)}\right)$
$=P\left(W^{(0)}=w^{(0)}\right)P\left(W^{(1)}=w^{(1)}\mid W^{(0)}=w^{(0)}\right)\cdots P\left(W^{(T)}=w^{(T)}\mid W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)}\right)$
$=P\left(W^{(0)}=1-w^{(0)}\right)P\left(W^{(1)}=1-w^{(1)}\mid W^{(0)}=1-w^{(0)}\right)\cdots$
$\hskip 15.0ptP\left(W^{(T)}=1-w^{(T)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)}\right)$
$=P\left(W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)},W^{(T)}=1-w^{(T)}\right),$
We then sum over all possible $\left(w^{(0)},\ldots,w^{(T-1)}\right)$ and obtain $P(W=w^{(T)})=P(W=1-w^{(T)})$. As a result, $P\left(W_{i}=1\right)=P\left(W_{i}=0\right)=1/2$. And we have
$E[\widehat{\tau}(W)]$ $=E\left[\frac{2}{n}\sum_{i=1}^{n}W_{i}Y_{i}(1)-\frac{2}{n}\sum_{i=1}^{n}\left(1-W_{i}\right)Y_{i}(0)\right]$
$=\frac{2}{n}\sum_{i=1}^{n}Y_{i}(1)P\left(W_{i}=1\right)-\frac{2}{n}\sum_{i=1}^{n}Y_{i}(0)P\left(W_{i}=0\right)$

### v2

@@ 5.2 Structured Randomized Experiments
This section considers the three types of structured experiments described in Section 4. We simulate the covariates and potential outcomes in the same way as in Section 5.1, except that we set $Y_{i}(1)=Y_{i}(0)+0.15\sqrt{\textnormal{Var}[Y_{i}(0)]}$ under the alternative to ensure the power of different methods remains within a moderate range. For sequential experiments, we set the number of groups as $K=2$ and consider two group sizes, $n_{k}\in\{100,500\}$. We use the strategy proposed by [45] to determine $a_{k}$. Specifically, we set $(s_{1},s_{2})=(239,761)$ when $p=50$ and $(s_{1},s_{2})=(264,736)$ when $p=250$, where $a_{k}=\frac{n_{k}}{n_{[k]}}q_{k}$ and $q_{k}$ is the $\frac{1}{s_{k}}$ quantile of a non-central chi-square distribution with $p$ degrees of freedom and a non-central parameter $\frac{n_{[k-1]}}{n_{k}}M_{[k-1]}$. For stratified experiments, we set the stratum number as $K=2$ and consider two stratum sizes, $n_{k}\in\{100,500\}$. For cluster randomized experiments, we set the cluster size as $n_{k}=2$ and consider two options for the number of clusters, $K\in\{100,500\}$. We set $p_{a}=P(\chi_{p}^{2}\leq a)=10^{-3}$ for both stratified randomized experiments and cluster randomized experiments and set $p=n/4$ for all settings in this section.
Table 2 shows the results when $n=1000$. The complete results are provided in Appendix D. For clarity, we use a unified name to refer to different method variants. Since PSRR is not extended to stratified experiments and cluster experiments, it is excluded from the two schemes. The overall findings are consistent with the results in Section 5.1. BRAIN substantially reduces the computational burden of rerandomization, especially for high-dimensional datasets and in stratified randomized experiments and cluster randomized experiments, where ARSRR is the only method available to perform rerandomization. BRAIN maintains the statistical properties of rerandomization in all the settings and significantly improves statistical inference compared to CR.
Table 2: Statistical and computational performance of different methods in structured randomized experiments. SD: standard deviation. CP: coverage probability. Len: length.
@@
Bias SD Size Power CP Len
@@
BRAIN 3.3 131 10.1 82 89.9 436 3.9
6 A Clinical Trial Example
In addition to the simulation studies, we analyze a real dataset from a phase 1 clinical trial [2] to illustrate the advantages of BRAIN. The goal of this trial was to evaluate the interactions between intravenous methamphetamine and oral reserpine. In this trial, 20 participants were randomly assigned to the treatment group, and 10 participants were assigned to the control group. We aim to balance three covariates: age, weight, and pre-treatment heart rate, which are significantly unbalanced in the original study, to better infer the treatment effects on post-treatment heart rate.
We set $a$ as the 0.001 quantile of $\chi_{3}^{2}$ and generate 10,000 acceptable assignments using each method. ARSRR takes approximately 2.3 minutes to complete the task, and PSRR takes more than half an hour. In contrast, BRAIN requires only 12 seconds, which demonstrates that BRAIN accelerates the rerandomization process significantly. Based on the original dataset, we simulate semi-synthetic data and examine the statistical performance of each method in Appendix D. The results are consistent with our simulation studies, where BRAIN preserves the desired statistical properties of rerandomization.
@@ 7 Conclusion
In this work, we propose a novel BRAIN method to perform rerandomization efficiently in various types of randomized experiments. We derive the unbiasedness and a lower bound for the variance reduction of the difference-in-means estimator under BRAIN. Our empirical results demonstrate that BRAIN maintains the appealing statistical properties of rerandomization and significantly reduces the computational cost compared to existing methods.
Limitations This work does not address randomized experiments with tiers of covariates. Rerandomization for such experiments requires searching for assignments that simultaneously balance the covariates within each tier. In future work, we plan to extend our method to this setting using multi-objective optimization techniques. Another limitation is that it is unclear whether other integer programming techniques (e.g., branch and bound) can be adapted to develop better rerandomization algorithms. We leave this as our future research direction.
@@
In Appendix A, we provide the pseudocode of the three BRAIN variants. In Appendix B, we state and prove all the theoretical results. In Appendix C, we introduce the Fisher randomization test and randomization-based confidence intervals. In Appendix D, we show the simulation results not included in the main text.
Appendix A Algorithmic Description of BRAIN Variants
A.1 Algorithmic Description of SeqBRAIN
The pseudocode of SeqBRAIN is presented in Algorithm 2. As mentioned in the main text, SeqBRAIN performs BRAIN in each experiment stage sequentially while fixing the assignments obtained in previous stages.
A.2 Algorithmic Description of StratBRAIN
The pseudocode of StratBRAIN is presented in Algorithm 3. StratBRAIN differs from BRAIN in that there are linear constraints within each stratum, $\sum_{i\in\mathcal{I}_{k}}W_{i}=n_{tk}$. Therefore, StratBRAIN only exchanges units in the same stratum to ensure that the aforementioned linear constraints are always satisfied. StratBRAIN does not loop over the strata sequentially. Instead, after selecting the pairs of units to examine from each stratum, StratBRAIN aggregates the lists of pairs and randomly permutes the combined list. In this way, we ensure that all strata are equally treated, allowing for flexible exploration of the neighborhood without a predetermined ordering of the strata or the units.
A.3 Algorithmic Description of ClustBRAIN
The pseudocode of ClustBRAIN is presented in Algorithm 4. ClustBRAIN treats each cluster as a whole and uses the clusters as the operational units in the search process. In other words, ClustBRAIN essentially conducts BRAIN at the cluster level rather than the unit level. Recall that we denote the assignment status of the clusters as a $K$-dimensional vector $U$ and all units’ assignment status as a $n$-dimensional vector $W$. We have the following constraint imposed by the cluster structure of the experiment. $W_{i}=U_{k}$ if the $k$-th cluster includes the $i$-th unit. Since $W$ is determined once we generate $U$, we denote $W(U)$ as the mapping from $U$ to $W$ in Algorithm 4.
@@ A.3 Algorithmic Description of ClustBRAIN
Algorithm 2 Sequential BRAIN
Input: Covariates $X$, index sets of units in each stratum $(\mathcal{I}_{1},\ldots,\mathcal{I}_{K})$, number of units to assign to the two arms in each stratum $(n_{t1},\ldots,n_{tK})$, $(n_{c1},\ldots,n_{cK})$, threshold $a$, hyperparameter $(L_{1},\ldots,L_{K})$, $(S_{1},\ldots,S_{K})$, $L_{\textnormal{total}}=\sum_{k}L_{k}$, $S_{\textnormal{total}}=\sum_{k}S_{k}$.
@@
       end for
      Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{L_{\textnormal{total}}})$ with random ordering;
       Aggregate $\cup_{1\leq k\leq K,1\leq l\leq L_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{L_{\textnormal{total}}})$. The ordering of $(Z_{1},\ldots,Z_{L_{\textnormal{total}}})$ is consistent with that of $(O_{1},\ldots,O_{L_{\textnormal{{total}}}})$;
       Set $\widetilde{S}=S_{\textnormal{total}}$;
       for $l=1,\ldots,L_{\textnormal{total}}$ do
@@
             end for
            Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{O_{kl}\}$ and denote the union as $(O_{1},\ldots,O_{S_{\textnormal{total}}})$;
             Aggregate $\cup_{1\leq k\leq K,1\leq l\leq S_{k}}\{Z_{kl}\}$ and denote the union as $(Z_{1},\ldots,Z_{S_{\textnormal{total}}})$;
             for $s=1,\ldots,S_{\textnormal{total}}$ do
@@
Algorithm 3 StratBRAIN
@@
Algorithm 4 ClustBRAIN
@@ Appendix B Additional Theoretical Results and Proofs
BRAIN, StratBRAIN, and ClustBRAIN all assign treatment in an offline manner. In contrast, SeqBRAIN performs online allocation and needs to be handled differently. We will first establish the theoretical properties of BRAIN, StratBRAIN, and ClustBRAIN, and then obtain the results of SeqBRAIN.
We have proposed the following conditions in the main text, which are related to the offline methods.
Condition 1.
@@
Similar to Condition 3, we also consider the following equal-group-size conditions for stratified randomized experiments and cluster randomized experiments, respectively.
Condition 4.
$n_{t_{k}}=n_{c_{k}}=\frac{n_{k}}{2}$, $k=1,\ldots,K$.
Condition 5.
@@
We point out three crucial properties of a rerandomization algorithm.
Property 1.
(symmetric feasible set) $\mathcal{F}=1-\mathcal{F}$.
Property 2.
(uniform initialization) For any $w\in\mathcal{F}$,
$\mathbb{P}\left(W^{(0)}=w\right)=\frac{1}{|\mathcal{F}|}.$
Property 3.
$\mathbb{P}\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)=u_{t}(M(w^{(0)}),\ldots,M(w^{(t)})),$
Any rerandomization algorithm with these properties can yield unbiased treatment effect estimators with reduced variances.
Theorem 3.
If $W$ is generated by a rerandomization algorithm satisfying Properties 1-3, then $E[\widehat{\tau}]=\tau$.
Theorem 4.
Under Conditions 1 and 2, if $W$ is generated by a rerandomization algorithm satisfying Properties 1-3, then
$\frac{\textnormal{V}_{\textnormal{CR}}-\textnormal{Var}[\widehat{\tau}]}{\textnormal{V}_{\textnormal{CR}}}\geq\left(1-\frac{a}{p}\right)R^{2},$
where $R^{2}=\frac{n\beta^{T}S_{XX}\beta}{n_{t}n_{c}\textnormal{V}_{\textnormal{CR}}}$.
We shall verify that BRAIN, StratBRAIN, and ClustBRAIN satisfy these properties.
Proposition 1.
BRAIN satisfies Properties 2 and 3. Under Condition 3, BRAIN satisfies Property 1.
Proposition 2.
StratBRAIN satisfies Properties 2 and 3. Under Condition 4, StratBRAIN satisfies Property 1.
Proposition 3.
ClustBRAIN satisfies Properties 2 and 3. Under Condition 5, ClustBRAIN satisfies Property 1.
Consequently, their unbiasedness and variance reduction results (including Theorems 1 and 2) are immediately evident as corollaries of Theorems 3 and 4.
In fact, the statistical guarantees of CR, ARSRR, and PSRR can also be established by verifying that they satisfy Properties 1-3.
We will establish the statistical properties of SeqBRAIN by extending the conditions and properties in the offline setting to the online setting.
@@ B.1 Proof of Theorem 3
$\mathbb{P}\left(W^{(0)}=w^{(0)}\right)=\mathbb{P}\left(W^{(0)}=1-w^{(0)}\right),$
@@
$\hskip 15.0pt\mathbb{P}\left(W^{(t)}=w^{(t)}\mid W^{(0)}=w^{(0)},\ldots,W^{(t-1)}=w^{(t-1)}\right)$
@@
$=\mathbb{P}\left(W^{(t)}=1-w^{(t)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(t-1)}=1-w^{(t-1)}\right),$
@@
$\hskip 15.0pt\mathbb{P}\left(W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)},W^{(T)}=w^{(T)}\right)$
$=\mathbb{P}\left(W^{(0)}=w^{(0)}\right)\mathbb{P}\left(W^{(1)}=w^{(1)}\mid W^{(0)}=w^{(0)}\right)\cdots\mathbb{P}\left(W^{(T)}=w^{(T)}\mid W^{(0)}=w^{(0)},\ldots,W^{(T-1)}=w^{(T-1)}\right)$
$=\mathbb{P}\left(W^{(0)}=1-w^{(0)}\right)\mathbb{P}\left(W^{(1)}=1-w^{(1)}\mid W^{(0)}=1-w^{(0)}\right)\cdots$
$\hskip 15.0pt\mathbb{P}\left(W^{(T)}=1-w^{(T)}\mid W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)}\right)$
$=\mathbb{P}\left(W^{(0)}=1-w^{(0)},\ldots,W^{(T-1)}=1-w^{(T-1)},W^{(T)}=1-w^{(T)}\right),$
We then sum over all possible $\left(w^{(0)},\ldots,w^{(T-1)}\right)$ and obtain $\mathbb{P}(W=w^{(T)})=\mathbb{P}(W=1-w^{(T)})$. As a result, $\mathbb{P}\left(W_{i}=1\right)=\mathbb{P}\left(W_{i}=0\right)=1/2$. And we have
$\mathbb{E}[\widehat{\tau}(W)]$ $=\mathbb{E}\left[\frac{2}{n}\sum_{i=1}^{n}W_{i}Y_{i}(1)-\frac{2}{n}\sum_{i=1}^{n}\left(1-W_{i}\right)Y_{i}(0)\right]$
$=\frac{2}{n}\sum_{i=1}^{n}Y_{i}(1)\mathbb{P}\left(W_{i}=1\right)-\frac{2}{n}\sum_{i=1}^{n}Y_{i}(0)\mathbb{P}\left(W_{i}=0\right)$